A canoeist travels 60 miles at a certain speed. The canoeist travels 18 miles on the second part of the trip at a speed of 5 mph slower. The total time for the trip is 3 hours, What was the speed on each part of the trip.
I tried this and just cant get it, would love to see your step by step solution.
Thankyou

Answer by scott8148(6628)

time equals distance divided by rate

(60 / x) + (18 / (x - 5)) = 3

multiplying by x(x - 5) ___ 60x - 300 + 18x = 3(x^2 - 5x)

0 = x^2 - 31x + 100

use quadratic formula to find x


this canoe must have a motor on it...